OpenStudy (anonymous):

If 6 ∫ f(x) dx 2 ≤ 6 ∫g(x) dx 2 , then f(x) ≤ g(x) for 2 ≤ x ≤ 6 . True False

OpenStudy (anonymous):

This is slightly better, however, still extremely hard to understand I'm not quite getting the question. @__assassin__ do you understand this?

OpenStudy (anonymous):

im looking at it but its hard to read the way you typed it

OpenStudy (anonymous):

I agree. This is very hard to understand.

OpenStudy (anonymous):

\[\int\limits_{2}^{6}f(x)dx \le \int\limits_{2}^{5}g(x)dx\]

OpenStudy (anonymous):

wait \[\int\limits_{2}^{6}g(x)dx\]

OpenStudy (anonymous):

I say it is false. Here's a counterexample. Even though one of the functions has a larger signed area, it can still get smaller than the other function in the interval.